A 200W-200V lamp is connected to 250V mains it power consumption is
 

Grade:12

1 Answers

Khimraj
5 years ago
power P = V2/R
R = 200*200/200 = 200
new power P = V2/R = 250*250/200 = 312.5 W
Hope it clears..................................................